OpenStudy (anonymous):

y= -5x+ 8 ...need answer and how to get answer

OpenStudy (angela210793):

wht to do with this? O.o

OpenStudy (anonymous):

it's a linear equation. i need to find x and y

OpenStudy (anonymous):

...you can type the equation in the calculator to find the value of x's and y's.

OpenStudy (anonymous):

or you can put any number in x... for example y=-5(1)+8 y= 3 so the x would be 1 and y would 3... (if that's what you need)

OpenStudy (anonymous):

i think for definite answer ..........u must have another equation to solve it...........

OpenStudy (anonymous):

well it depends on the question and what he's tring to ask... :) so I really don't know if he's just trying to find all the domain and range...

OpenStudy (anonymous):

I was looking for the x and y intercepts of this equation so i can graph a line

OpenStudy (anonymous):

to find y-int plug in 0 for x and to find x-int plug in 0 for y

OpenStudy (anonymous):

to find y intercept put x=0 as y=-5(0) +8 and y=8 and to find x intercept put y=0 u will get 0=-5x+8 5x=8 in this way x=8/5

OpenStudy (anonymous):

s0 value of x and y will be (8/5,8)

OpenStudy (anonymous):

thank you